# Auslander-Reiten translations in the monomorphism categories of exact categories

### v1

@@ Abstract.
Let $\Lambda$ be a finite dimensional algebra. Let ${\mathcal{C}}$ be a functorially finite exact subcategory of $\Lambda$-mod with enough projective and injective objects and ${\mathcal{S}}({\mathcal{C}})$ be its monomorphism category. It turns out that the category ${\mathcal{S}}({\mathcal{C}})$ has almost split sequences. We show an explicit formula for the Auslander-Reiten translation in ${\mathcal{S}}({\mathcal{C}})$. Furthermore, if ${\mathcal{C}}$ is a stably $d$-Calabi-Yau Frobenius category, we calculate objects under powers of Auslander-Reiten translation in the triangulated category $\overline{{\mathcal{S}}({\mathcal{C}})}$.
@@ 1. Introduction
Throughout the paper, $\Lambda$ stands for a finite dimensional algebra over a field $k$ and $\Lambda\textendash\operatorname{mod}$ the category of finitely generated left $\Lambda$-modules. Denote by $D=\operatorname{Hom}_{k}(-,k)$ the $k$-dual functor.
@@ 1.1. Background
The classification of subobjects is a fundamental problem in mathematics. In [Birk], G. Birkhoff proposed the problem of classifying all subgroups of abelian $p$-groups by matrices. This question is related to representations of partial order sets, such as [Ar, Si1, Si2]. C. M. Ringel and M. Schmidmeier [RS1, RS2, RS3] revisited this subject using a modern point of view by studying the representation type and the Auslander-Reiten theory of the submodule category for artin algebras $k[x]/(x^{n})$, for $n>1$. In [RS2], they developed the Auslander-Reiten theory in the submodule category for an artin algebra and showed that the Ausladner-Reiten translation in the submodule category ${\mathcal{S}}(\Lambda)$ can be computed by the Auslander-Reiten translation within the module category.
@@
In this case, it is well-known that ${\mathcal{S}}({\rm Gproj}(\Lambda))$ is a functorially finite subcategory of $T_{2}(\Lambda)\textendash\operatorname{mod}$, and hence it has almost split sequences. In [LZ3], the authors showed a similar result about the existence of almost split sequences in monomorphism categories.
@@
If ${\mathcal{C}}$ is contravariantly finite resolving in subcategory in $\Lambda\textendash\operatorname{mod}$, then so is ${\mathcal{S}}({\mathcal{C}})$ in $T_{2}(\Lambda)\textendash\operatorname{mod}$ and hence it has almost split sequences.
@@ 1.2. Main results
Motivated by these results above, we will focus on studying the existence of almost split sequences in the monomorphism category of an exact category and their Auslander-Reiten translations in this paper. We will show an explicit formula for the Auslander-Reiten translation in the monomorphism category ${\mathcal{S}}({\mathcal{C}})$ and the epimorphism category ${\mathcal{F}}({\mathcal{C}})$.
@@
Let $\Lambda$ be a finite dimensional algebra and ${\mathcal{C}}$ be a functorially finite exact subcategory of $\Lambda\textendash\operatorname{mod}$ with enough projective and injective objects. Then the followings hold:
@@
    Some special almost split sequences in ${\mathcal{H}}({\mathcal{C}})$, ${\mathcal{S}}({\mathcal{C}})$ and ${\mathcal{F}}({\mathcal{C}})$ are classified in Lemma 3.9 and Lemma 6.3 respectively. The others are so-called cw-exact i.e. degreewise split exact (see Definition 3.1).
@@
    RSS-equivalence. Let $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C$ be a conflation in an exact category ${\mathcal{C}}$. We denote by $\operatorname{Coker}f=g$ and $\operatorname{Ker}g=f$. In fact, this induces a pair of equivalence functors $\operatorname{\bf Cok}:{\mathcal{S}}({\mathcal{C}})\to{\mathcal{F}}({\mathcal{C}})$ and $\operatorname{\bf Ker}:{\mathcal{F}}({\mathcal{C}})\to{\mathcal{S}}({\mathcal{C}})$, which are called the RSS-equivalences. Furthermore, if ${\mathcal{C}}$ has enough projective objects, they induce functors on the stable categories $\operatorname{\bf Cok}:\underline{{\mathcal{S}}({\mathcal{C}})}\to\underline{{\mathcal{F}}({\mathcal{C}})}$ and $\operatorname{\bf Ker}:\underline{{\mathcal{F}}({\mathcal{C}})}\to\underline{{\mathcal{S}}({\mathcal{C}})}$.
@@
    $\operatorname{\bf Mimo}$ and epivalence. The notion of minimal monomorphism was introduced in [RS2]. For any morphism $f:A\to B$ in a module category $\Lambda\textendash\operatorname{mod}$, the minimal monomorphism $\operatorname{\bf Mimo}f=[f,e]^{T}:A\to B\oplus I\operatorname{Ker}f$, where $e$ is an extension of the injective envelope $i:\operatorname{Ker}f\to I\operatorname{Ker}f$. It was shown that $\operatorname{\bf Mimo}f$ is unique up to isomorphism, independent on the choice of $e$ and it is a minimal right ${\mathcal{S}}(\Lambda)$-approximation of $f$. In [BM], the authors pointed out that $\operatorname{\bf Mimo}$ is a functor ${\mathcal{H}}(\Lambda)\to\overline{{\mathcal{S}}(\Lambda)}$. In [LS, GKKP2], it was shown that right ${\mathcal{S}}({\mathcal{C}})$-approximations $\mathfrak{R}f\to f$ give rise to a functor $\mathfrak{R}:{\mathcal{H}}(\Lambda)\to\overline{{\mathcal{S}}(\Lambda)}$. In this paper, we generalize this notion for a Krull-Schmidt exact category ${\mathcal{C}}$ with enough injective objects and denote the image of right ${\mathcal{S}}({\mathcal{C}})$-approximations of $f$ in the stable subcategory $\overline{{\mathcal{S}}({\mathcal{C}})}$ by $\operatorname{\bf Mimo}f$. Hence, $\operatorname{\bf Mimo}:{\mathcal{H}}({\mathcal{C}})\to\overline{{\mathcal{S}}({\mathcal{C}})}$ is a functor.
    Another important functor is the epivalence $\pi:\overline{{\mathcal{S}}({\mathcal{C}})}\to{\mathcal{H}}(\overline{\mathcal{C}})$ (see Definition 2.5, Theorem 5.6), which sends a monomorphism $f$ to its image $\overline{f}$ in the stable category $\overline{\mathcal{C}}$ and $\pi^{\prime}:\underline{{\mathcal{F}}({\mathcal{C}})}\to{\mathcal{H}}(\underline{{\mathcal{C}}})$ which sends an epimorphism $f$ to its image $\underline{f}$ in $\underline{{\mathcal{C}}}$. Since an epivalence induces a bijection of indecomposable objects between categories, we can define a map $\operatorname{\bf Mimo}$ on the objects of ${\mathcal{H}}(\overline{\mathcal{C}})$ which agrees with the inverse of $\pi$ (Remark 5.9).
@@
    Functoriality. We clarify the notations in these two Auslander-Reiten formulas. In general, if a Krull-Schmidt exact category ${\mathcal{C}}$ with enough projective and injective objects has almost split sequences, then the Auslander-Reiten translation is a functor $\tau_{\mathcal{C}}:\underline{{\mathcal{C}}}\to\overline{\mathcal{C}}$. However, our formulas for $\tau_{\mathcal{S}}$ and $\tau_{\mathcal{F}}$ only hold for objects $f$ in $\underline{{\mathcal{S}}({\mathcal{C}})}$ or $\underline{{\mathcal{F}}({\mathcal{C}})}$, but not for morphisms. Well-definedness of the formulas can be checked directly from the following compositions:
@@
    We briefly explain the strategy of our proof here. First of all, up to some reductions it suffices to prove the formula $\tau_{\mathcal{F}}(f)\cong\operatorname{\bf Cok}\operatorname{\bf Mimo}\tau_{\mathcal{C}}(\underline{f})$ holds for epimorphisms $f$ which is the end term of a cw-exact almost split sequence (see Section 6.3). We use the stable functor ${\mathcal{F}}({\mathcal{C}})\to{\mathcal{H}}(\underline{{\mathcal{C}}})$ ($f\mapsto\underline{f}$) to transfer cw-exact almost split sequences in ${\mathcal{F}}({\mathcal{C}})$ to cw-exact almost split sequences in ${\mathcal{H}}(\underline{{\mathcal{C}}})$ (Lemma 6.5). And then we can use the functor $\Theta:{\mathcal{H}}(\underline{{\mathcal{C}}})\to\underline{{\mathcal{C}}}\textendash\operatorname{mod}$ to transfer cw-exact almost split sequences in ${\mathcal{H}}(\underline{{\mathcal{C}}})$ to almost split sequences in $\underline{{\mathcal{C}}}\textendash\operatorname{mod}$. Finally, the problem boils down to computing the Auslander-Reiten translation on $\underline{{\mathcal{C}}}\textendash\operatorname{mod}$, which has been investigated by [AR74].
@@
    Generalizations and relevant results. In this paper, we focus on the “submodule categories” as its importance of being a prototype of the general (separated) monomorphism categories which have been studied intensively in [LZ1, LZ2, ZX]. Meanwhile, further investigations and generalizations are expected. Possible generalizations are two-folded. On one hand, in contrast of our assumption that ${\mathcal{C}}$ is a subcategory of some ambient module category and has enough projective and injective objects, one may consider the Auslander-Reiten theory of ${\mathcal{S}}({\mathcal{C}})$ for a $\operatorname{Hom}$-finite Krull-Schmidt exact category ${\mathcal{C}}$ having almost split sequences (e.g. the category of coherent sheaves ${\rm coh}\mathbb{P}^{1}$). For such categories ${\mathcal{C}}$, it is beyond our knowledge so far to show that ${\mathcal{S}}({\mathcal{C}})$ also has almost split sequences. On the other hand, one may consider the Auslander-Reiten theory for separated monomorphism categories ${\rm smon}(Q,{\mathcal{C}})$ of an exact category ${\mathcal{C}}$ over an acyclic quiver $Q$. In [XZZ], the Auslander-Reiten translation formula was given for $Q$ an $\mathbb{A}_{n}$ type quiver with linear orientation and ${\mathcal{C}}=\Lambda\textendash\operatorname{mod}$ for some artin algebra $\Lambda$. In [GKKP3], the authors explicitly described the Auslander-Reiten translation for the monomorphism category of dualisable pro-species.
@@
Let ${\mathcal{C}}$ be a functorially finite stably $d$-Calabi-Yau Frobenius category of $\Lambda\textendash\operatorname{mod}$. Let $\langle 1\rangle$ be the suspension functor of the stable category $\overline{{\mathcal{S}}({\mathcal{C}})}$. The Auslander-Reiten translation $\tau_{\mathcal{S}}:\overline{{\mathcal{S}}({\mathcal{C}})}\to\overline{{\mathcal{S}}({\mathcal{C}})}$ satisfies $\tau_{\mathcal{S}}^{6}f\cong f\langle 6d-4\rangle$ for every object $f$ in $\overline{{\mathcal{S}}({\mathcal{C}})}$.
@@ 2.1. Exact categories and (co)stable categories
An exact structure ${\mathcal{E}}$ on an additive category ${\mathcal{C}}$ is a fixed class of composable morphism pairs $A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C$, where $i$ is a kernel of $p$ and $p$ is a cokernel of $i$, satisfying certain axioms, see e.g. [B, Definition 2.1]. An exact category $({\mathcal{C}},{\mathcal{E}})$ is a pair of an additive category ${\mathcal{C}}$ and an exact structure ${\mathcal{E}}$ on it. An element $A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C$ of ${\mathcal{E}}$ is called a conflation or short exact sequence, where $i$ is called an inflaction or admissible monomorphism and $p$ is called a deflation or admissible epimorphism. Later on, when there is no confusion, we simply call ${\mathcal{C}}$ an exact category, omitting the exact structure ${\mathcal{E}}$ and call inflations (resp. deflations and conflations) monomorphisms (resp. epimorphisms and exact sequences). A conflation is denoted by $0\to A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C\to 0$.
Let $({\mathcal{C}},{\mathcal{E}})$ be an exact category. A full subcategory ${\mathcal{D}}\subseteq{\mathcal{C}}$ is called an exact subcategory of $({\mathcal{C}},{\mathcal{E}})$ if ${\mathcal{D}}$ is closed under extensions and equipped with an exact structure which contains $A\to B\to C\in{\mathcal{E}}$, where $A,B,C\in{\mathcal{D}}$.
In an exact category $({\mathcal{C}},{\mathcal{E}})$, an object $P$ is called projective if for any deflation $p:B\to C$, any morphism $u:P\to C$ can factor through $p$. An epimorphism $p:P(X)\to X$ is called a projective cover, if $P(X)$ is a projective object in ${\mathcal{C}}$ and any endomorphism $\phi\in\operatorname{End}_{\mathcal{C}}(P(x))$ satisfying $\phi=p\circ\phi$ is an automorphism. We say that an exact category $({\mathcal{C}},{\mathcal{E}})$ has enough projectives if for any object $C$ in ${\mathcal{C}}$, there exists a deflation $p:P\to C$ with $P$ a projective object. Notions with respect to injective objects and injective envelopes can be defined dually.
@@
In some literature, projective objects are also called $\operatorname{Ext}$-projective and exact subcategory are called fully exact subcategory
Let ${\mathcal{C}}$ be an exact category. Denote by $\underline{{\mathcal{C}}}$ the stable category, whose objects are the same as ${\mathcal{C}}$ and morphism sets are $\operatorname{Hom}_{\underline{{\mathcal{C}}}}(X,Y)=\operatorname{Hom}_{{\mathcal{C}}}(X,Y)/{\mathcal{P}}(X,Y)$, where ${\mathcal{P}}(X,Y)$ denotes the set of morphisms $X\to Y$ which can factor through a projective object. The costable category $\overline{\mathcal{C}}$ is defined dually. Morphisms in $\underline{{\mathcal{C}}}$ (resp. $\overline{\mathcal{C}}$) are usually denoted by $\underline{f}$ (resp.$\overline{f}$). Define $Q_{\mathcal{C}}:{\mathcal{C}}\to\underline{{\mathcal{C}}}$ the stable functor which sends morphisms $f\mapsto\underline{f}$. The costable functor $Q^{\mathcal{C}}:{\mathcal{C}}\to\overline{\mathcal{C}}$ is defined dually. A Frobenius category is an exact category $({\mathcal{C}},{\mathcal{E}})$ with enough projectives and enough injectives, moreover, the projectives coincide with the injectives.

### v2

@@ Abstract.
Let $\mathcal{V}$} be a dualizing $k$-variety over a field $k$ and $\mathcal{C}$} be a functorially finite exact subcategory of ${\mathcal{V}}\textendash\operatorname{mod}$ having enough $\operatorname{Ext}$-projective and $\operatorname{Ext}$-injective objects. Let $\mathcal{H}(\mathcal{C})$ (resp. $\mathcal{S}(\mathcal{C})$ and $\mathcal{F}(\mathcal{C})$) be its morphism (resp. monomorphism and epimorphism) category. It turns out that these categories have almost split sequences. We show an explicit formula for the Auslander-Reiten translation in $\mathcal{S}(\mathcal{C})$. Furthermore, if $\mathcal{C}$ is a stably $d$-Calabi-Yau Frobenius category, we calculate objects under powers of Auslander-Reiten translation in the triangulated category $\overline{\mathcal{S}(\mathcal{C})}$.
@@ 1. Introduction
Throughout the paper, $k$ is a field and $\Lambda$ is a finite dimensional algebra over $k$. Denote by $D=\operatorname{Hom}_{k}(-,k)$ the $k$-dual functor.
@@ 1.1. Background
The classification of subobjects is a fundamental problem in mathematics. In [Birk], G. Birkhoff proposed the problem of classifying all subgroups of abelian $p$-groups by matrices. This question is related to representations of partial order sets, such as [Ar, Si1, Si2]. C. M. Ringel and M. Schmidmeier [RS1, RS2, RS3] revisited this subject using a modern point of view by studying the representation type and the Auslander-Reiten theory of the submodule category for artin algebras $k[x]/(x^{n})$, for $n>1$. In [RS2], they developed the Auslander-Reiten theory in the submodule category for an artin algebra and showed that the Auslander-Reiten translation in the submodule category ${\mathcal{S}}(\Lambda)$ can be computed by the Auslander-Reiten translation within the module category.
@@
In this case, it is well-known that ${\mathcal{S}}({\rm Gproj}(\Lambda))$ is a functorially finite subcategory of $T_{2}(\Lambda)\textendash\operatorname{mod}$, and hence it has almost split sequences. In [LZ3], the authors showed a similar result about the existence of almost split sequences in the monomorphism categories of complete hereditary cotorsion classes.
@@
If $\mathcal{C}$} is a contravariantly finite resolving subcategory in $\Lambda\textendash\operatorname{mod}$, then so is ${\mathcal{S}}({\mathcal{C}})$ in $T_{2}(\Lambda)\textendash\operatorname{mod}$ and hence it has almost split sequences.
@@ 1.2. Main results
Motivated by these results above, we will focus on studying the existence of almost split sequences in the monomorphism category of an exact category $\mathcal{C}$} and their Auslander-Reiten translations in this paper. We will show an explicit formula for the Auslander-Reiten translation in the monomorphism category ${\mathcal{S}}({\mathcal{C}})$ and the epimorphism category ${\mathcal{F}}({\mathcal{C}})$.
@@
Let $\mathcal{V}$} be a dualizing $k$-variety and $\mathcal{C}$} be a functorially finite exact subcategory of ${\mathcal{V}}\textendash\operatorname{mod}$ having enough $\operatorname{Ext}$-projective and $\operatorname{Ext}$-injective objects. Then the followings hold:
@@
    Some special almost split sequences in ${\mathcal{H}}({\mathcal{C}})$, ${\mathcal{S}}({\mathcal{C}})$ and ${\mathcal{F}}({\mathcal{C}})$ are classified in Lemma 3.8. The others are so-called cw-exact i.e. degreewise split exact (see Definition 3.2).
@@
    RSS-equivalence. Let $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C$ be a conflation in an exact category $\mathcal{C}$}. We denote by $\operatorname{Coker}f=g$ and $\operatorname{Ker}g=f$. In fact, this induces a pair of equivalence functors $\operatorname{\bf Cok}:{\mathcal{S}}({\mathcal{C}})\to{\mathcal{F}}({\mathcal{C}})$ and $\operatorname{\bf Ker}:{\mathcal{F}}({\mathcal{C}})\to{\mathcal{S}}({\mathcal{C}})$, which are called the RSS-equivalences. Furthermore, if $\mathcal{C}$} has enough $\operatorname{Ext}$-projective objects, they induce functors on the stable categories $\operatorname{\bf Cok}:\underline{{\mathcal{S}}({\mathcal{C}})}\to\underline{{\mathcal{F}}({\mathcal{C}})}$ and $\operatorname{\bf Ker}:\underline{{\mathcal{F}}({\mathcal{C}})}\to\underline{{\mathcal{S}}({\mathcal{C}})}$.
@@
    $\operatorname{\bf Mimo}$ and epivalence. The notion of minimal monomorphism was introduced in [RS2]. For any morphism $f:A\to B$ in a module category $\Lambda\textendash\operatorname{mod}$, the minimal monomorphism $\operatorname{\bf Mimo}f=[f,e]^{T}:A\to B\oplus I\operatorname{Ker}f$, where $e$ is an extension of the injective envelope $i:\operatorname{Ker}f\to I\operatorname{Ker}f$ by the inclusion $\operatorname{Ker}f\hookrightarrow A$. It was shown that $\operatorname{\bf Mimo}f$ is unique up to isomorphism, independent on the choice of $e$ and it is a minimal right ${\mathcal{S}}(\Lambda)$-approximation of $f$. In [BM], the authors pointed out that $\operatorname{\bf Mimo}$ is a functor ${\mathcal{H}}(\Lambda)\to\overline{{\mathcal{S}}(\Lambda)}$. In [LS, GKKP2], it was shown that right ${\mathcal{S}}(\Lambda)$-approximations $\mathfrak{R}f\to f$ give rise to a functor $\mathfrak{R}:{\mathcal{H}}(\Lambda)\to\overline{{\mathcal{S}}(\Lambda)}$. In this paper, we generalize this notion for a Krull-Schmidt exact category $\mathcal{C}$} with enough $\operatorname{Ext}$-injective objects and denote the image of right ${\mathcal{S}}({\mathcal{C}})$-approximations of $f$ in the stable subcategory $\overline{{\mathcal{S}}({\mathcal{C}})}$ by $\operatorname{\bf Mimo}f$. Hence, $\operatorname{\bf Mimo}:{\mathcal{H}}({\mathcal{C}})\to\overline{{\mathcal{S}}({\mathcal{C}})}$ is a functor.
    Another important functor is the epivalence $\pi:\overline{{\mathcal{S}}({\mathcal{C}})}\to{\mathcal{H}}(\overline{{\mathcal{C}}})$ (see Definition 2.7, Theorem 5.6), which sends an inflation $f$ to its image $\overline{f}$ in the stable category $\overline{{\mathcal{C}}}$ and $\pi^{\prime}:\underline{{\mathcal{F}}({\mathcal{C}})}\to{\mathcal{H}}(\underline{{\mathcal{C}}})$ which sends a deflation $f$ to its image $\underline{f}$ in $\underline{{\mathcal{C}}}$. Since an epivalence induces a bijection of indecomposable objects between categories, it turns out that the inverse image of an indecomposable object $\overline{f}$ in ${\mathcal{H}}(\overline{{\mathcal{C}}})$ under the functor $\pi$ agrees with $\operatorname{\bf Mimo}f$. Hence we denote $\operatorname{\bf Mimo}\overline{f}:=\operatorname{\bf Mimo}f$ as an object in $\overline{{\mathcal{S}}({\mathcal{C}})}$ (Remark 5.9).
@@
    Functoriality. We clarify the notations in these two Auslander-Reiten formulas. In general, if a Krull-Schmidt exact category $\mathcal{C}$} with enough $\operatorname{Ext}$-projective and $\operatorname{Ext}$-injective objects has almost split sequences, then the Auslander-Reiten translation is a functor $\tau_{\mathcal{C}}:\underline{{\mathcal{C}}}\to\overline{{\mathcal{C}}}$. However, our formulas for $\tau_{\mathcal{S}}$ and $\tau_{\mathcal{F}}$ only hold for objects $f$ in $\underline{{\mathcal{S}}({\mathcal{C}})}$ or $\underline{{\mathcal{F}}({\mathcal{C}})}$, but not for morphisms $(\alpha,\beta):f\to g$ in $\underline{{\mathcal{S}}({\mathcal{C}})}$ or $\underline{{\mathcal{F}}({\mathcal{C}})}$. Well-definedness of the formulas can be checked directly from the following compositions:
@@
    We briefly explain the strategy of our proof here. First of all, up to some reductions it suffices to prove the formula $\tau_{\mathcal{F}}(f)\cong\operatorname{\bf Cok}\operatorname{\bf Mimo}\tau_{\mathcal{C}}(\underline{f})$ holds for deflations $f$ which is the end term of a cw-exact almost split sequence (see Section 6.2). We use the stable functor ${\mathcal{F}}({\mathcal{C}})\to{\mathcal{H}}(\underline{{\mathcal{C}}})$ ($f\mapsto\underline{f}$) to transfer cw-exact almost split sequences in ${\mathcal{F}}({\mathcal{C}})$ to cw-exact almost split sequences in ${\mathcal{H}}(\underline{{\mathcal{C}}})$ (Lemma 6.3). And then we can use the functor $\Theta:{\mathcal{H}}(\underline{{\mathcal{C}}})\to\underline{{\mathcal{C}}}\textendash\operatorname{mod}$ to transfer cw-exact almost split sequences in ${\mathcal{H}}(\underline{{\mathcal{C}}})$ to almost split sequences in $\underline{{\mathcal{C}}}\textendash\operatorname{mod}$. Finally, the problem boils down to computing the Auslander-Reiten translation on $\underline{{\mathcal{C}}}\textendash\operatorname{mod}$, which has been investigated by [AR74].
@@
    Generalizations and relevant results. In this paper, we focus on the “submodule categories” as its importance of being a prototype of the general (separated) monomorphism categories which have been studied intensively in [LZ1, LZ2, ZX]. Meanwhile, further investigations and generalizations are expected. Possible generalizations are two-folded. On one hand, in contrast of our assumption that $\mathcal{C}$} is an exact subcategory of an abelian category and has enough $\operatorname{Ext}$-projective and $\operatorname{Ext}$-injective objects, one may consider the Auslander-Reiten theory of ${\mathcal{S}}({\mathcal{C}})$ for a $\operatorname{Hom}$-finite Krull-Schmidt exact category $\mathcal{C}$} having almost split sequences, but without enough $\operatorname{Ext}$-projective or $\operatorname{Ext}$-injective objects (e.g. the category of coherent sheaves ${\rm coh}\mathbb{P}^{1}$). For such categories $\mathcal{C}$}, it is beyond our knowledge so far to show that ${\mathcal{S}}({\mathcal{C}})$ also has almost split sequences. On the other hand, one may consider the Auslander-Reiten theory for separated monomorphism categories ${\rm smon}(Q,{\mathcal{C}})$ of an exact category $\mathcal{C}$} over an acyclic quiver $Q$. In [XZZ], the Auslander-Reiten translation formula was given for an $\mathbb{A}_{n}$ type quiver $Q$ with linear orientation and ${\mathcal{C}}=\Lambda\textendash\operatorname{mod}$ for some artin algebra $\Lambda$. In [GKKP3], the authors explicitly described the Auslander-Reiten translation for the monomorphism category of dualisable pro-species.
@@
Let $\mathcal{V}$} be a dualizing $k$-variety and $\mathcal{C}$} be a functorially finite stably $d$-Calabi-Yau Frobenius category of ${\mathcal{V}}\textendash\operatorname{mod}$. Let $\langle 1\rangle$ be the suspension functor of the stable category $\overline{{\mathcal{S}}({\mathcal{C}})}$. The Auslander-Reiten translation $\tau_{\mathcal{S}}:\overline{{\mathcal{S}}({\mathcal{C}})}\to\overline{{\mathcal{S}}({\mathcal{C}})}$ satisfies $\tau_{\mathcal{S}}^{6}f\cong f\langle 6d-4\rangle$ for every object $f$ in $\overline{{\mathcal{S}}({\mathcal{C}})}$.
@@ 2.1. Exact categories and (co)stable categories
An exact structure $\mathcal{E}$} on an additive category $\mathcal{C}$} is a fixed class of composable morphism pairs $A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C$, where $i$ is a kernel of $p$ and $p$ is a cokernel of $i$, satisfying certain axioms, see e.g. [B, Definition 2.1]. An exact category $({\mathcal{C}},{\mathcal{E}})$ is a pair of an additive category $\mathcal{C}$} and an exact structure $\mathcal{E}$} on it. An element $A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C$ of $\mathcal{E}$} is called a conflation or short exact sequence, where $i$ is called an inflation or admissible monomorphism and $p$ is called a deflation or admissible epimorphism. Later on, when there is no confusion, we simply call $\mathcal{C}$} an exact category, omitting the exact structure $\mathcal{E}$}. A conflation is usually denoted by $0\to A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C\to 0$.
Example 2.1.
For any additive category $\mathcal{C}$}, denote by ${\mathcal{E}}^{\rm split}$ the exact structure containing only split exact sequences in $\mathcal{C}$}. Then ${\mathcal{C}}^{\rm split}:=({\mathcal{C}},{\mathcal{E}}^{\rm split})$ is an exact category. In fact, ${\mathcal{E}}^{\rm split}$ is the smallest exact structure on $\mathcal{C}$}.
Let $({\mathcal{C}},{\mathcal{E}})$ and $({\mathcal{C}}^{\prime},{\mathcal{E}}^{\prime})$ be exact categories. We say an additive functor $F:{\mathcal{C}}\to{\mathcal{C}}^{\prime}$ preserves exact sequences if $FA\stackrel{{\scriptstyle Ff}}{{\to}}FB\stackrel{{\scriptstyle Fg}}{{\to}}FC\in{\mathcal{E}}^{\prime}$ whenever $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C\in{\mathcal{E}}.$ We say $F$ reflects exact sequences, if $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C\in{\mathcal{E}}$ whenever $FA\stackrel{{\scriptstyle Ff}}{{\to}}FB\stackrel{{\scriptstyle Fg}}{{\to}}FC\in{\mathcal{E}}^{\prime}$. An exact functor is an additive functor which preserves exact sequences. An exact equivalence is an additive equivalence functor $F:({\mathcal{C}},{\mathcal{E}})\to({\mathcal{C}}^{\prime},{\mathcal{E}}^{\prime})$, which preserves and reflects exact sequences.
Example 2.2.
Let $({\mathcal{C}},{\mathcal{E}})$ be an exact category and $\mathcal{E}$} contain at least one non-split exact sequences. Then the identity functor $1:({\mathcal{C}},{\mathcal{E}}^{\rm split})\to({\mathcal{C}},{\mathcal{E}})$ is an exact functor, but not an exact equivalence.
Let $({\mathcal{C}},{\mathcal{E}})$ be an exact category. A full subcategory ${\mathcal{D}}\subseteq{\mathcal{C}}$ is called an exact subcategory of $({\mathcal{C}},{\mathcal{E}})$ if $\mathcal{D}$} is closed under extensions and equipped with an exact structure which contains $A\to B\to C\in{\mathcal{E}}$, where $A,B,C\in{\mathcal{D}}$. In some literature, exact subcategory are called fully exact subcategory.
In an exact category $({\mathcal{C}},{\mathcal{E}})$, conflations $0\to A\to B\to C\to 0$ form an abelian group $\operatorname{Ext}^{1}(C,A)$ under Baer sum, which is also an $\operatorname{End}(A)$-$\operatorname{End}(C)$-bimodule with the actions given by pushouts and pullbakcs respectively. A morphism $f:X\to Y$ is called projectively trivial if the induced map $\operatorname{Ext}^{1}(f,M):\operatorname{Ext}^{1}(Y,M)\to\operatorname{Ext}^{1}(X,M)$ is zero for any object $M\in{\mathcal{C}}$. An object $P$ is called $\operatorname{Ext}$-projective if the identity $1_{P}$ is projectively trivial. We say $\mathcal{C}$} has enough $\operatorname{Ext}$-projective objects if for each $X\in{\mathcal{C}}$, there is a deflation $f:P\to X$ for some $\operatorname{Ext}$-projective object $P$. Notions with respect to injectively trivial morphisms and $\operatorname{Ext}$-injective objects can be defined dually.
@@
When $\mathcal{C}$} has enough $\operatorname{Ext}$-projective (injective) objects, a morphism $f$ is projectively (injectively) trivial if and only if it factors through an $\operatorname{Ext}$-projective (injective) object. When $\mathcal{C}$} is an abelian category and $\mathcal{E}$} consists of all the short exact sequences in $\mathcal{C}$}, then $\operatorname{Ext}$-projective (injective) objects are simply projective and injective objects in $\mathcal{C}$}.
Let $\mathcal{C}$} be an exact category. Denote by $\underline{{\mathcal{C}}}$ the stable category, whose objects are the same as $\mathcal{C}$} and morphism sets are $\operatorname{Hom}_{\underline{{\mathcal{C}}}}(X,Y)=\operatorname{Hom}_{{\mathcal{C}}}(X,Y)/{\mathcal{P}}(X,Y)$, where ${\mathcal{P}}(X,Y)$ denotes the set of projectively trivial morphisms from $X$ to $Y$. The costable category $\overline{{\mathcal{C}}}$ is defined dually. Given a morphism $f\in{\mathcal{C}}$, denote by $\underline{f}$ (resp. $\overline{f}$) the class of morphisms represented by $f$ in $\underline{{\mathcal{C}}}$ (resp. $\overline{{\mathcal{C}}}$). Define $Q_{\mathcal{C}}:{\mathcal{C}}\to\underline{{\mathcal{C}}}$ the stable functor which sends morphisms $f\mapsto\underline{f}$. The costable functor $Q^{\mathcal{C}}:{\mathcal{C}}\to\overline{{\mathcal{C}}}$ is defined dually. A Frobenius category is an exact category $({\mathcal{C}},{\mathcal{E}})$ with enough $\operatorname{Ext}$-projectives and $\operatorname{Ext}$-injectives, moreover, the $\operatorname{Ext}$-projectives coincide with the $\operatorname{Ext}$-injectives.